the speed of light in a vacuum is 3.00 x 10^8 m/s. if the reflective index of diamond is 2.42 what speed does a light travel as it passes through a diamond?


1
Expert's answer
2021-05-21T17:05:56-0400
"v=\dfrac{c}{n}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{2.42}=1.24\cdot10^8\ \dfrac{m}{s}."